What is the probability that someone consumed less than 20 gallons of bottled​ water?


1
Expert's answer
2021-09-27T18:14:12-0400

The annual per capita consumption of bottled water was 31.9 gallons. Assume that the per capita consumption of bottled water is approximately normally distributed with a mean of 31.9 and a standard deviation of 10 gallons.

What is the probability that someone consumed less than 20 gallons of bottled​ water?


"P(x<20)=P(\n\n\n\n\\frac{x\u2013\u03bc}{\\sigma}\n\n\u200b<\n\n\n\n\\frac{20\u201331.9}{10}\n\n\u200b)=P(z<\u22121.19)=0.1170"